# Convergent Privacy Framework with Contractive GNN Layers for Multi-hop Aggregations

### v1

@@ 4.3. Privacy Allocation Module
Building on Section 4.2, we integrate CAM and PAM to design the Perturbed Contractive Message Passing (PCMP) (Algorithm 1). By leveraging contractiveness, PCMP ensures bounded privacy loss for long $K$-hop graphs, eliminating the linear growth of privacy loss with $K$. A subsequent projection step enforces Lipschitz constraints, maintaining consistent scaling across hops. After $K$ iterations, PCMP outputs private feature matrices $\hat{\mathbf{X}}^{(K)}$, which are passed to the classification module.
Table 2. Comparison between Private Message Passing GNNs. “Utility” summarizes the results of $\mathsf{GAP}$ and $\mathsf{CARIBOU}$ in Table 3. “Chains” is the chain-structured dataset (Gu et al., 2020) and “Standard” represents other real-world graph datasets.
Framework Contractiveness Node Embeddings Calibrated Noise ($\sigma$) Privacy Loss Utility (Chains) Utility (Standard) GAP $\times$ Intermediate & final $\propto\sqrt{K}$ Unbounded $\mathsf{CARIBOU}$checkmark Final $\propto\sqrt{\min(K,\frac{1-C_{\mathsf{L}}^{K}}{1+C_{\mathsf{L}}^{K}}\frac{1+C_{\mathsf{L}}}{1-C_{\mathsf{L}}})}$ Bounded
@@
4$\triangleright$ Calculate the Required Noise Calibration (from PAM).
@@
6:     Calculate $\Delta(\mathrm{MP})$ through Equation 6
@@
8:     Calculate $\Delta(\mathrm{MP})$ through Equation 8
@@
12$\triangleright$ Perturbed Message Passing with Contractiveness (from CAM).
@@
15:     $\triangleright$ Contractive graph layer: compute the node embeddings.
16:     ${{\bm{X}}}^{(k+1)}\leftarrow{{\bm{X}}}^{(k+1)}+\mathcal{N}(\mu,(\Delta(\mathsf{CGL}))^{2}\sigma^{2})$triangleright Perturbation.
17:     ${{\bm{X}}}^{(k+1)}\leftarrow\Pi_{\mathcal{K}}({{\bm{X}}}^{(k)})$triangleright Projection constrained by norm $1$.
18:end for
19:
20:Return: ${\mathbf{X}}^{(K)}$
Privacy Budget Allocation.
The privacy budget is distributed across $K$ hops to ensure the total privacy loss adheres to the specified $\epsilon$ and $\delta$. The noise scale is calibrated based on the sensitivity of graph operations and the desired DP guarantees, using a noise allocation mechanism (NAM) that limits noise accumulation under Lipschitz constraints. The maximum allowable noise calibration (see Corollary 10) is constrained by $K^{\prime}=\min(K,(1+C_{\mathsf{L}})/(1-C_{\mathsf{L}}))$.
@@
4.4. Privacy Auditing Module
Message passing mechanisms integrate graph structures by recursively aggregating and transforming node features from neighbors. Membership Inference Attacks (MIA) on graph examine the vulnerability of message passing to infer whether specific nodes or edges were part of a GNN’s training set (Olatunji et al., 2021b). The adversaries exploit black-box access to the GNN, querying it with selected data and analyzing outputs (e.g., class probabilities) to infer membership. To align theoretical privacy guarantees with practical deployments, we propose an empirical auditing module. This module simulates an adversary to evaluate GNN privacy before deployment. Extending Carlini et al. (Carlini et al., 2022)’s MIA framework to graphs, we define a graph-specific privacy auditing game (Definition 1) and implement two real-world attacks (Wu et al., 2022a; Olatunji et al., 2021b) for privacy verification.
@@ 4.4. Privacy Auditing Module
  2. (2)
    The challenger flips a bit $b$, and if $b=0$, samples a fresh challenge point from the distribution $(x,y)\leftarrow\mathbb{G}$ (such that $(x,y)\notin G$). Otherwise, the challenger selects a point from the training set $(x,y)\overset{\$}{\leftarrow}G$.
  3. (3)
@@ 4.5. Putting it Together
To improve classification accuracy, $\mathsf{CARIBOU}$ adopts a head MLP architecture proposed by Sajadmanesh et al. (Sajadmanesh et al., 2023) as the CM. This design ensures that the CM effectively combines the information from both feature sets, enabling robust node classification. Furthermore, the CM guarantees efficient training by leveraging the graph-agnostic features ${\bm{X}}^{(0)}$, ensuring a lower-bound performance even in scenarios where graph topology is unavailable.
5. Convergent Privacy Analysis
In this section, we present a convergent privacy analysis for perturbed message-passing GNNs with respect to the number of hops. We begin by reviewing the standard privacy analysis for a one-layer perturbed message-passing GNN, and then illustrate that the privacy cost grows linearly with the number of layers under standard composition theorems. We then shift our focus to a convergent privacy analysis for perturbed GNNs whose message-passing layers are contractive. In particular, we draw upon the framework of contractive noisy iteration (CNI) from Bok et al. (2024), recasting the multi-layer perturbed GNN as a CNI process. Our analysis reveals that, under hidden intermediate states and contractive message-passing layers, the privacy cost converges as the number of hops increases. Finally, we specialize this result to our proposed $\mathsf{CARIBOU}$. We show that $\mathsf{CARIBOU}$’s message-passing operation is contractive, derive its sensitivity, and thereby establish concrete bounds on both edge-level and node-level differential privacy for arbitrarily many hops. All proofs are deferred to the Appendix C.
5.1. Preliminaries: Perturbed Message Passing
Recall that each layer of a perturbed message-passing GNN updates the node features following Equation 2, where ${\bm{Z}}\sim\mathcal{N}\bigl{(}0,\sigma^{2}\mathrm{I}_{d}\bigr{)}$ is Gaussian noise added to obscure the true output of $\mathrm{MP}_{{\bm{G}}},$ and $\Pi_{\mathcal{K}}$ projects onto some convex feasible set ${\mathcal{K}}.$ As shown in Proposition 1, for a single layer, the privacy cost depends only on $\Delta(\mathrm{MP})$ and $\sigma.$ However, by the composition theorem of RDP (Mironov, 2017), stacking $K$ such layers yields a cost that scales linearly in $K$ (Corollary 2).
Proposition 0.
[…]
Corollary 0.
[…]
This linear growth of the privacy budget is common in multi-hop settings, such as (Sajadmanesh et al., 2023), but is troublesome for deep GNNs, especially when many message-passing steps are needed for long-range dependencies. Excessively large privacy loss often undermines the model’s utility.
5.2. Contractive Noisy Iteration and Convergent Privacy
Many GNNs, such as GCN, use "mean-type" aggregation, mixing a node’s representation with that of its neighbors. Intuitively, iterative mixing could "amplify" privacy, but existing analyses yield only linear compositions. Our key observation is that perturbed message passing in a contractive GNN layer behaves analogously to noisy gradient descent or noisy iterative maps (Chourasia et al., 2021; Altschuler and Talwar, 2022; Altschuler et al., 2024), where recent work has demonstrated privacy amplification via iteration.
Below, we introduce the framework of contractive noisy iteration (CNI) from Bok et al. (2024), and the meta theorem proved by Bok et al. (2024) that provides a tight privacy guarantee for CNI processes.
Definition 0 (Contractive Noisy Iteration (CNI)(Bok et al., 2024, Definition 3.1)).
@@ 5.2. Contractive Noisy Iteration and Convergent Privacy
(5) ${\bm{X}}^{(k+1)}=\Pi_{\mathcal{K}}\!\bigl{(}\,\phi_{k+1}\bigl{(}{\bm{X}}^{(k)}\bigr{)}+{\bm{Z}}^{(k)}\bigr{)},$
@@
A special case of CNI considered in (Bok et al., 2024) is the noisy gradient descent, where the contractive functions are the gradient update steps for some fixed loss function $f$, and the noise distribution $\xi_{k}$ is Gaussian. This situation is similar to the perturbed message passing mechanism in a contractive GNN layer, where the contractive function $\phi_{k}\equiv\mathrm{MP}_{G}$ is the message passing operation for a fixed graph ${\bm{G}}$.
The complete privacy analysis of CNI processes in Bok et al. (2024) involves concepts of trade-off functions (Definition 1) and Gaussian differential privacy (GDP, Definition 2), which we briefly review in Appendix B. GDP can be converted to the more familiar Rényi Differential Privacy (RDP) framework:
Lemma 0 (GDP Implies RDP (Bok et al., 2024, Lemma A.4)).
[…]
With this connection established, we now state the key meta theorem from Bok et al. (2024) that analyzes the CNI process and provides a tight privacy guarantee.
Theorem 5 (Meta Theorem on CNI (Bok et al., 2024, Theorem C.5)).
@@
The theorem above slightly generalizes the original result from Bok et al. (2024) by relaxing the Lipschitz condition to require $\gamma<1$ only for $k\geq 2$ rather than for all iterations. This relaxation is critical for analyzing our $\mathsf{CARIBOU}$ architecture, where the first message-passing layer includes the residual term $\beta{\bm{X}}^{(0)}$, potentially making it non-contractive while subsequent layers remain contractive. The proof extends the original argument by carefully tracking the influence of the first layer on the privacy guarantee. For completeness, we provide the full proof in Appendix C.
@@
Theorem 6 ((Bok et al., 2024, Theorem 4.2) adapted for GNNs).
@@
[…]
@@
5.3. Edge- and Node-Level Privacy of $\mathsf{CARIBOU}$
@@
[…]
where $10\leq C_{\mathsf{L}}<$ 1, and hyper-parameters, $\alpha_{1},\alpha_{2},\beta>0,\alpha_{1}+\alpha_{2}=1$. In order to establish the privacy guarantees of $\mathsf{CARIBOU}$, we need to determine the sensitivity and the Lipschitz constant of the message passing layer $\mathsf{CGL}$.
@@
(6) $\Delta_{e}(\mathsf{CGL})=\sqrt{2}C_{\mathsf{L}}\alpha_{1}\Big{(}$ $\frac{1}{(D_{\min}+1)(D_{\min}+2)}$
$+\frac{C(D_{\min})}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}\sqrt{D_{\min}+1}}\Big{)},$
@@
(7) $C(D_{\min})=\begin{cases}\frac{D_{\min}}{\sqrt{D_{\min}+1}}-\frac{D_{\min}}{\sqrt{D_{\min}+2}}&D_{\min}>3\\ (\frac{3}{\sqrt{4}}-\frac{3}{\sqrt{5}})&1\leq D_{\min}\leq 3\\ \end{cases}$
@@
(8) $\Delta_{n}(\mathsf{CGL})=1+\quad\alpha_{2}C_{\mathsf{L}}\frac{2|V|}{|V|+1}$
$+\alpha_{1}C_{\mathsf{L}}\left(\frac{\sqrt{D_{\max}}}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})\sqrt{D_{\max}}}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}}\right)$
@@
Then as a consequence of the contractive result in Proposition 7 and the sensitivity results in Theorem 8 and Theorem 9, we can apply Theorem 6 to derive the following privacy guarantee for the message-passing layer $\mathsf{CGL}$ of $\mathsf{CARIBOU}$.

### v2

@@ IV-B Privacy Allocation Module
Building on Section IV-A, we integrate CAM and PAM to design the Perturbed Contractive Message Passing (PCMP) (Algorithm 1). By leveraging contractiveness, PCMP ensures bounded privacy loss for long $K$-hop graphs, eliminating the linear growth of privacy loss with $K$. A subsequent projection step enforces Lipschitz constraints, maintaining consistent scaling across hops. After $K$ iterations, PCMP outputs private feature matrices $\hat{\mathbf{X}}^{(K)}$, which are passed to the classification module.
@@
[…]
@@
6:  Calculate $\Delta(\mathsf{CGL})$ through Equation 6
@@
8:  Calculate $\Delta(\mathsf{CGL})$ through Equation 8
@@
[…]
@@
15:  $\triangleright$ Contractive graph layer: compute node embeddings.
16:  ${{\bm{X}}}^{(k+1)}\leftarrow{{\bm{X}}}^{(k+1)}+\mathcal{N}(\mu,(\Delta(\mathsf{CGL}))^{2}\sigma^{2})$
17:  $\triangleright$ DP Perturbation.
18:  ${{\bm{X}}}^{(k+1)}\leftarrow\Pi_{\mathcal{K}}({{\bm{X}}}^{(k)})$ $\triangleright$ Projection with norm 11.
19:end for
20:
21:Return: ${\mathbf{X}}^{(K)}$
Privacy Budget Allocation. The privacy budget is distributed across $K$ hops to ensure the total privacy loss adheres to the specified $\epsilon$ and $\delta$. The noise scale is calibrated based on the sensitivity of graph operations and the desired DP guarantees, using a noise allocation mechanism (NAM) that limits noise accumulation under Lipschitz constraints. The maximum allowable noise calibration (see Corollary 1) is constrained by $K^{\prime}=\min(K,(1+C_{\mathsf{L}})/(1-C_{\mathsf{L}}))$.
@@
IV-C Privacy Auditing Module
Message passing mechanisms integrate graph structures by recursively aggregating and transforming node features from neighbors. Membership Inference Attacks (MIA) on graph examine the vulnerability of message passing to infer whether specific nodes or edges were part of a GNN’s training set [tpsisa/OlatunjiNK21]. The adversaries exploit black-box access to the GNN, querying it with selected data and analyzing outputs (e.g., class probabilities) to infer membership. To align theoretical privacy guarantees with practical deployments, we propose an empirical auditing module. This module simulates an adversary to evaluate GNN privacy before deployment. Extending Carlini et al. [sp/CarliniCN0TT22]’s MIA framework to graphs, we define a graph-specific privacy auditing game (Definition 7) and implement two real-world attacks [sp/Wulinkteller22, tpsisa/OlatunjiNK21] for privacy verification.
@@ IV-C Privacy Auditing Module
  2. 2.
    The challenger flips a bit $b$, and if $b=0$, samples a fresh challenge point from the distribution $(x,y)\leftarrow\mathbb{G}$ (such that $(x,y)\notin G$). Otherwise, the challenger selects a point from the training set $(x,y)\overset{\mathdollar}{\leftarrow}G$.
  3. 3.
@@ IV-D Putting it Together
To improve classification accuracy, $\mathsf{CARIBOU}$ adopts a head MLP architecture proposed by Sajadmanesh et al. [uss/sajadmanesh2023gap] as the CM. This design ensures that the CM effectively combines the information from both feature sets, enabling robust node classification. Furthermore, the CM guarantees efficient training by leveraging the graph-agnostic features ${\bm{X}}^{(0)}$, ensuring a lower-bound performance even in scenarios where graph topology is unavailable.
V Convergent Privacy Analysis
In this section, we present a convergent privacy analysis for perturbed message-passing GNNs with respect to the number of hops. In Appendix A-A, we review the standard privacy analysis for a one-layer perturbed message-passing GNN, and then illustrate that the privacy cost grows linearly with the number of layers under standard composition theorems. We then shift our focus to a convergent privacy analysis for perturbed GNNs whose message-passing layers are contractive. In particular, we draw upon the framework of contractive noisy iteration (CNI) from [icml/BokShifted24], recasting the multi-layer perturbed GNN as a CNI process. Our analysis reveals that, under hidden intermediate states and contractive message-passing layers, the privacy cost converges as the number of hops increases. Finally, we specialize this result to our proposed $\mathsf{CARIBOU}$. We show that $\mathsf{CARIBOU}$’s message-passing operation is contractive, derive its sensitivity, and thereby establish concrete bounds on both edge-level and node-level differential privacy for arbitrarily many hops. All proofs are deferred to the Appendices A-B,A-C.
V-A Contractive Noisy Iteration and Convergent Privacy
Many GNNs, such as GCN, use "mean-type" aggregation, mixing a node’s representation with that of its neighbors. Intuitively, iterative mixing could "amplify" privacy, but existing analyses yield only linear compositions. Our key observation is that perturbed message passing in a contractive GNN layer behaves analogously to noisy gradient descent or noisy iterative maps [nips/ChourasiaYS21, nips/AltschulerT22, siamcomp/AltschulerBT24], where recent work has demonstrated privacy amplification via iteration.
Below, we introduce the framework of contractive noisy iteration (CNI) from [icml/BokShifted24], and the meta theorem proved by [icml/BokShifted24] that provides a tight privacy guarantee for CNI processes.
Definition 8 (Contractive Noisy Iteration (CNI)[icml/BokShifted24, Definition 3.1]).
@@ V-A Contractive Noisy Iteration and Convergent Privacy
${\bm{X}}^{(k+1)}=\Pi_{\mathcal{K}}\!\bigl(\,\phi_{k+1}\bigl({\bm{X}}^{(k)}\bigr)+{\bm{Z}}^{(k)}\bigr),$ (5)
@@
A special case of CNI considered in [icml/BokShifted24] is the noisy gradient descent, where the contractive functions are the gradient update steps for some fixed loss function $f$, and the noise distribution $\xi_{k}$ is Gaussian. This situation is similar to the perturbed message passing mechanism in a contractive GNN layer, where the contractive function $\phi_{k}\equiv\mathrm{MP}_{G}$ is the message passing operation for a fixed graph $\bm{G}$}.
The complete privacy analysis of CNI processes in [icml/BokShifted24] involves concepts of trade-off functions (Definition 9) and Gaussian differential privacy (GDP, Definition 10), which we briefly review in Appendix A. GDP can be converted to the more familiar Rényi Differential Privacy (RDP) framework:
Lemma 1 (GDP Implies RDP [icml/BokShifted24, Lemma A.4]).
If a mechanism is $\mu\text{-GDP},$ then it satisfies $\left(\alpha,\frac{1}{2}\,\alpha\,\mu^{2}\right)\text{-RDP}$ for all $\alpha>1.$
With this connection established, we now state the key meta theorem from [icml/BokShifted24] that analyzes the CNI process and provides a tight privacy guarantee.
Theorem 2 (Meta Theorem on CNI [icml/BokShifted24, Theorem C.5]).
@@
The theorem above slightly generalizes the original result from [icml/BokShifted24] by relaxing the Lipschitz condition to require $\gamma<1$ only for $k\geq 2$ rather than for all iterations. This relaxation is critical for analyzing our $\mathsf{CARIBOU}$ architecture, where the first message-passing layer includes the residual term $\beta{\bm{X}}^{(0)}$, potentially making it non-contractive while subsequent layers remain contractive. The proof extends the original argument by carefully tracking the influence of the first layer on the privacy guarantee. For completeness, we provide the full proof in Appendices A-B,A-C.
@@
Theorem 3 ([icml/BokShifted24, Theorem 4.2] adapted for GNNs).
@@
$\left(\alpha,\frac{1}{2}\alpha\frac{\Delta^{2}(\mathrm{MP})}{\sigma^{2}}\frac{1-\gamma^{K}}{1+\gamma^{K}}\frac{1+\gamma}{1-\gamma}\right)\text{-RDP},$
which is equivalent to $\left(\frac{1}{2}\alpha\frac{\Delta^{2}(\mathrm{MP})}{\sigma^{2}}\frac{1-\gamma^{K}}{1+\gamma^{K}}\frac{1+\gamma}{1-\gamma}+\frac{\log(1/\delta)}{\alpha-1},\delta\right)\text{-DP}.$
@@
V-B Edge- and Node-Level Privacy of $\mathsf{CARIBOU}$
@@
$\text{CGL:}\quad\quad{\bm{X}}^{(k+1)}=C_{\mathsf{L}}(\alpha_{1}\hat{{\bm{A}}}{\bm{X}}^{(k)}+\alpha_{2}\mathsf{Mean}({\bm{X}}^{(k)}))+\beta{\bm{X}}^{(0)},$
where $0\leq C_{\mathsf{L}}<1$, and hyper-parameters, $\alpha_{1},\alpha_{2},\beta>0,\alpha_{1}+\alpha_{2}=1$. In order to establish the privacy guarantees of $\mathsf{CARIBOU}$, we need to determine the sensitivity and the Lipschitz constant of the message passing layer $\mathsf{CGL}$.
@@
$\Delta_{e}(\mathsf{CGL})=\sqrt{2}C_{\mathsf{L}}\alpha_{1}\Big($ $\frac{1}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})}{\sqrt{D_{\min}+1}}$ (6)
$+\frac{1}{\sqrt{D_{\min}+2}\sqrt{D_{\min}+1}}\Big),$
@@
[…]
@@
$\Delta_{n}(\mathsf{CGL})=1+\quad\alpha_{2}C_{\mathsf{L}}\frac{2|V|}{|V|+1}+\alpha_{1}C_{\mathsf{L}}\left(\frac{\sqrt{D_{\max}}}{(D_{\min}+1)(D_{\min}+2)}\right$. (8)
$\qquad\left.+\frac{C(D_{\min})\sqrt{D_{\max}}}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}}\right)$
@@
Then as a consequence of the contractive result in Proposition 1 and the sensitivity results in Theorem 4 and Theorem 5, we can apply Theorem 3 to derive the following privacy guarantee (Corollary 1) for the message-passing layer $\mathsf{CGL}$ of $\mathsf{CARIBOU}$. Specifically, for Lipchitz constant $0.8$, a $10$-layer GNN realizes $(\alpha,{3.62\alpha\Delta^{2}}/{\sigma^{2}})$-RDP (versus $(\alpha,{5\alpha\Delta^{2}}/{\sigma^{2}})$-RDP for GAP), meanwhile $2$-layer GNN is $(\alpha,{0.99\alpha\Delta^{2}}/{\sigma^{2}})$-RDP (versus $(\alpha,{\alpha\Delta^{2}}/{\sigma^{2}})$-RDP for GAP). That is, more privacy costs are saved as the number of layer increases.
@@
$\left(\alpha,\frac{\alpha}{2}\frac{\Delta^{2}}{\sigma^{2}}\min\left\{K,\frac{1-C_{\mathsf{L}}^{K}}{1+C_{\mathsf{L}}^{K}}\frac{1+C_{\mathsf{L}}}{1-C_{\mathsf{L}}}\right\}\right)\text{-RDP},$
where $\Delta=\Delta_{e}(\mathsf{CGL})$ from Theorem 4 for edge-level privacy, or $\Delta=\Delta_{n}(\mathsf{CGL})$ from Theorem 5 for node-level privacy. In particular, as $K\to\infty$, it converges to $\left(\alpha,\frac{\alpha}{2}\frac{\Delta^{2}}{\sigma^{2}}\frac{1+C_{\mathsf{L}}}{1-C_{\mathsf{L}}}\right)\text{-RDP}.$ Also, the RDP guarantees imply the following DP guarantees:
$\left(\frac{\alpha}{2}\frac{\Delta^{2}}{\sigma^{2}}\min\left\{K,\frac{1-C_{\mathsf{L}}^{K}}{1+C_{\mathsf{L}}^{K}}\frac{1+C_{\mathsf{L}}}{1-C_{\mathsf{L}}}\right\}+\frac{\log(1/\delta)}{\alpha-1},\delta\right)\text{-DP}.$
Theorem 6 ($\mathsf{CARIBOU}$’s Privacy Guarantee).
For any $\alpha>1$, let DAE training and CM training satisfy $(\alpha,\epsilon_{\mathsf{DAE}}(\alpha))$-RDP and $(\alpha,\epsilon_{\mathsf{CM}}(\alpha))$-RDP, respectively. Then, for the maximum hop $K\geq 1$, privacy failure probability $0<\delta<1$, Lipschitz constant $0<C_{\mathsf{L}}<1$, and noise variance $\sigma^{2}$, $\mathsf{CARIBOU}$ satisfies $(\epsilon_{\mathsf{CARIBOU}},\delta)$-DP, where $\epsilon_{\mathsf{CARIBOU}}=\epsilon_{\mathsf{DAE}}(\alpha)+\epsilon_{\mathsf{CM}}(\alpha)+\frac{\alpha}{2}\frac{\Delta^{2}}{\sigma^{2}}\min\left\{K,\frac{1-C_{\mathsf{L}}^{K}}{1+C_{\mathsf{L}}^{K}}\frac{1+C_{\mathsf{L}}}{1-C_{\mathsf{L}}}\right\}+\frac{\log(1/\delta)}{\alpha-1}.$ Here, $\Delta=\Delta_{e}(\mathsf{CGL})$ from Theorem 4 for edge-level privacy, or $\Delta=\Delta_{n}(\mathsf{CGL})$ from Theorem 5 for node-level privacy.
TABLE III: Comparison of Classification Accuracy for EDP and NDP. The best accuracy and the second-best accuracy are highlighted, respectively. The symbol